An ice-making machine inside a refrigerator operates in a Carnot cycle. It takes heat from liquid water at 0.0∘C and rejects heat to a room at a temperature of 24.5 ∘C. Suppose that liquid water with a mass of 84.4 kg at 0.0∘C is converted to ice at the same temperature. Take the heat of fusion for water to be Lf = 3.34×105 J/kg . Part A How much heat |QH| is rejected to the room? Express your answer in joules to four significant figures. Answer in J
Answer #1

The heat absorbed by the carnot engine when the ice freezes

Q=mLf =84.4*(3.34*105) =-2.82*107 J

QC = 2.82*107J

For Carnot cycle

QC/QH = TC/TH

=>QH=QC*(TH/TC) =(2.82*106)(273+24.5/273)

QH = 3.07*107 J